# Temperature Scaling Attack Disrupting Model Confidence in Federated Learning

### v1

@@ Appendix B Experimental Details
Table 4: Federated learning and attack configuration across dataset–model pairs (default settings unless stated otherwise).
B.1 Federated learning setup
@@
Table 5: Configuration used for case studies in Section 5.6.
Figure 17: Example client label distributions under Dirichlet partitioning for CIFAR-10 with different concentration parameters $\alpha$ (smaller $\alpha$ yields more heterogeneous splits).
@@
To model statistical heterogeneity, we partition each dataset across clients using a Dirichlet split [hsu2019measuring]. Specifically, for each class $c$, we draw a client-wise proportion vector $\boldsymbol{\pi}_{c}\sim\mathrm{Dirichlet}(\alpha)$ and allocate samples of class $c$ to clients according to $\boldsymbol{\pi}_{c}$. The concentration parameter $\alpha$ controls the non-IID severity: smaller $\alpha$ yields more skewed, label-dominant clients, while larger $\alpha$ approaches an approximately IID split. Figure 17 visualizes representative label distributions under different $\alpha$. As shown, the data distribution emulates both the quantity and label heterogeneity. Unless otherwise specified, we use $\alpha=1.0$, and we sweep $\alpha$ in our sensitivity experiments (Fig. 10).
@@
B.2 Centered Kernel Alignment
@@
$\mathrm{CKA}(X,Y)\triangleq\frac{\|\tilde{X}^{\top}\tilde{Y}\|_{F}^{2}}{\|\tilde{X}^{\top}\tilde{X}\|_{F}\;\|\tilde{Y}^{\top}\tilde{Y}\|_{F}}.$ (30)
@@
B.3 Defenses and post-hoc calibration
@@
$\bullet$ Robust aggregation defenses. FedAvg [mcmahan2017communication] is the standard baseline that aggregates client updates by data-size weights and performs no explicit adversary filtering. FLTrust [cao2020fltrust] assigns trust scores to client updates using a small server-held trusted dataset. Concretely, the server computes a reference update on the trusted data and reweights each client update by its cosine similarity to the reference direction (often with additional norm clipping), suppressing updates that are poorly aligned with the trusted gradient. FoolsGold [fung2018mitigating] targets sybil-style poisoning by downweighting clients whose updates are too similar to others over time. It maintains per-client similarity statistics and reduces the aggregation weight of clients that repeatedly submit highly correlated update directions, under the intuition that colluding attackers tend to produce clustered gradients. MultiKrum [blanchard2017machine] is a distance-based Byzantine-robust rule. For each client update, it computes the sum of distances to its closest neighbors and selects one or several updates with the smallest such score (i.e., those most consistent with the majority), then aggregates the selected set. This aims to filter out outliers that deviate strongly in update space.
Figure 18: Impact of calibration distortion on threshold-based verification in mobile healthcare triage in percentage change (%). Left: unnecessary verifications. Right: missed verifications, versus threshold $\gamma$.
Figure 19: Extended results for selective perception under temperature-induced miscalibration in DETR.
@@
Figure 20: Additional confidence histograms and reliability diagrams under different training temperatures $\tau$. Across both settings, varying $\tau$ consistently shifts the confidence distribution and reliability curve while leaving accuracy largely stable.
$\bullet$ Post-hoc calibration methods. For post-hoc calibration (Fig. 13), we evaluate temperature scaling [guo2017calibration], Platt scaling [platt1999probabilistic], isotonic regression [zadrozny2002transforming], and histogram binning [zadrozny2001obtaining]. All methods fit a mapping from an uncalibrated model score to a calibrated confidence using a held-out calibration set. Temperature scaling [guo2017calibration] fits a single scalar $T>0$ that rescales logits at inference, $\mathbf{p}=\mathrm{softmax}(\mathbf{z}/T)$, by minimizing negative log-likelihood on the calibration set. It preserves the predicted class (logit ordering) and only adjusts probability sharpness. Platt scaling [platt1999probabilistic] fits a parametric sigmoid mapping (typically on a scalar score such as a logit margin), $p(y{=}1\mid s)=\sigma(as+b)$, via maximum likelihood on the calibration set. It is a two-parameter, monotone calibrator. Isotonic regression [zadrozny2002transforming] fits a non-parametric monotone piecewise-constant function that maps scores to probabilities by minimizing squared error subject to monotonicity. It is flexible but can overfit when the calibration set is small. Histogram binning [zadrozny2001obtaining] partitions scores into bins and replaces each score by the empirical accuracy within its bin. This provides a simple non-parametric estimate of $P(\text{correct}\mid s)$, with a bias–variance trade-off controlled by the number of bins and calibration-set size.

### v2

@@ Appendix B Experimental Details
Table 4. Federated learning and attack configuration across dataset–model pairs (default settings unless stated otherwise).
B.1. Federated learning setup
@@
Table 5. Configuration used for case studies in Section 5.6.
Figure 17. Example client label distributions under Dirichlet partitioning for CIFAR-10 with different concentration parameters $\alpha$ (smaller $\alpha$ yields more heterogeneous splits).
@@
To model statistical heterogeneity, we partition each dataset across clients using a Dirichlet split [21]. Specifically, for each class $c$, we draw a client-wise proportion vector $\boldsymbol{\pi}_{c}\sim\mathrm{Dirichlet}(\alpha)$ and allocate samples of class $c$ to clients according to $\boldsymbol{\pi}_{c}$. The concentration parameter $\alpha$ controls the non-IID severity: smaller $\alpha$ yields more skewed, label-dominant clients, while larger $\alpha$ approaches an approximately IID split. Figure 17 visualizes representative label distributions under different $\alpha$. As shown, the data distribution emulates both the quantity and label heterogeneity. Unless otherwise specified, we use $\alpha=1.0$, and we sweep $\alpha$ in our sensitivity experiments (Fig. 10).
@@
B.2. Centered Kernel Alignment
@@
(30) $\mathrm{CKA}(X,Y)\triangleq\frac{\|\tilde{X}^{\top}\tilde{Y}\|_{F}^{2}}{\|\tilde{X}^{\top}\tilde{X}\|_{F}\;\|\tilde{Y}^{\top}\tilde{Y}\|_{F}}.$
@@
B.3. Defenses and post-hoc calibration
@@
$\bullet$ Robust aggregation defenses. FedAvg [36] is the standard baseline that aggregates client updates by data-size weights and performs no explicit adversary filtering. FLTrust [6] assigns trust scores to client updates using a small server-held trusted dataset. Concretely, the server computes a reference update on the trusted data and reweights each client update by its cosine similarity to the reference direction (often with additional norm clipping), suppressing updates that are poorly aligned with the trusted gradient. FoolsGold [12] targets sybil-style poisoning by downweighting clients whose updates are too similar to others over time. It maintains per-client similarity statistics and reduces the aggregation weight of clients that repeatedly submit highly correlated update directions, under the intuition that colluding attackers tend to produce clustered gradients. MultiKrum [5] is a distance-based Byzantine-robust rule. For each client update, it computes the sum of distances to its closest neighbors and selects one or several updates with the smallest such score (i.e., those most consistent with the majority), then aggregates the selected set. This aims to filter out outliers that deviate strongly in update space.
Figure 18. Impact of calibration distortion on threshold-based verification in mobile healthcare triage in percentage change (%). Left: unnecessary verifications. Right: missed verifications, versus threshold $\gamma$.
Figure 19. Extended results for selective perception under temperature-induced miscalibration in DETR.
@@
Figure 20. Additional confidence histograms and reliability diagrams under different training temperatures $\tau$. Across both settings, varying $\tau$ consistently shifts the confidence distribution and reliability curve while leaving accuracy largely stable.
$\bullet$ Post-hoc calibration methods. For post-hoc calibration (Fig. 13), we evaluate temperature scaling [17], Platt scaling [45], isotonic regression [51], and histogram binning [50]. All methods fit a mapping from an uncalibrated model score to a calibrated confidence using a held-out calibration set. Temperature scaling [17] fits a single scalar $T>0$ that rescales logits at inference, $\mathbf{p}=\mathrm{softmax}(\mathbf{z}/T)$, by minimizing negative log-likelihood on the calibration set. It preserves the predicted class (logit ordering) and only adjusts probability sharpness. Platt scaling [45] fits a parametric sigmoid mapping (typically on a scalar score such as a logit margin), $p(y{=}1\mid s)=\sigma(as+b)$, via maximum likelihood on the calibration set. It is a two-parameter, monotone calibrator. Isotonic regression [51] fits a non-parametric monotone piecewise-constant function that maps scores to probabilities by minimizing squared error subject to monotonicity. It is flexible but can overfit when the calibration set is small. Histogram binning [50] partitions scores into bins and replaces each score by the empirical accuracy within its bin. This provides a simple non-parametric estimate of $P(\text{correct}\mid s)$, with a bias–variance trade-off controlled by the number of bins and calibration-set size.